Define absolutely continuous function on [a,b] .suppose f is absolutely continuous, prove that |f| is absolutely continuous.


1
Expert's answer
2021-03-25T04:17:24-0400

A function "f:[a,b] \to \mathbb{R}" is absolutely continuous on "[a,b]" if for every "\epsilon >0 \, \exist \, \delta >0" such that whenever a finite sequence of pairwise disjoint subintervals "(x_k,y_k)" of "[a,b]" with "x_k,y_k \in [a,b]" satisfies "\sum_k (y_k -x_k) < \delta" then "\sum_k |f(y_k)-f(x_k) | < \epsilon"


Suppose that f is absolutely continuous. This implies that the condition of the definition above has been satisfied.

By triangle inequality, we have that

"||f(y_k)| - |f(x_k)|| \leq |f(y_k)-f(x_k)|"

"\implies \sum_k||f(y_k)| - |f(x_k)|| \leq \sum_k |f(y_k)-f(x_k)| < \epsilon \\
\therefore \sum_k||f(y_k)| - |f(x_k)|| < \epsilon"

Hence |f| is absolutely continuous.